OpenStudy (anonymous):

in the binomial theorem how the n(n-1)/2! came please explain

OpenStudy (anonymous):

there are a lot more terms than \(\frac{n(n-1)}{2}\)

OpenStudy (amistre64):

lots of expanding binomial powers

OpenStudy (amistre64):

a pattern develops, the coeffs relate to pascals triangle

OpenStudy (anonymous):

\(\frac{n(n-1)}{2}\) is the number of ways you can choose 2 items from a set of \(n\) commonly called "n choose 2" and written as \(\binom{n}{2}\) or \(_nC_2\)

OpenStudy (amistre64):

\[(a+b) = 1a+1b\] \[(a+b)^2 = 1a^2+2ab+1b^2\] \[(a+b)^3 = 1a^3+3a^2b+3ab^2+1b^3\] \[(a+b)^4 = 1,4,6,4,1\] \[(a+b)^5 = 1,5,10,10,5,1\] etc ...

OpenStudy (amistre64):

there is some induction method that then applies to form a proof is all

OpenStudy (anonymous):

you can also see it, if you have the patience, by actually computing this and not using the commutative law it is messy, but it becomes very clear why the coefficients are "n choose k"

OpenStudy (amistre64):

1 1 1 1 2 1 1 3 3 1 1 4 6 4 1 1 (1+4) (4+6) (6+4) (4+1) 1 etc ....

OpenStudy (anonymous):

for example \[(a+b)(a+b)(a+b)=(aa+ab+ba+bb)(a+b)\] \[=aaa+aba+baa+abb+aab+abb+bab+bbb\]

OpenStudy (amistre64):

messy indeed :)

OpenStudy (anonymous):

true, but it takes the mystery out of the \(\binom{n}{k}\) part right? how many terms look like \(a^2b\)? \(aab,aba, baa\) the number of ways you can choose 2 out of three slots to put the \(a\)

OpenStudy (anonymous):

all of this is obscured somewhat by the rather useful commutative law but if you were working non commutative system, things would be different